Hi!
I am having a problem setting up and solving this problem. I know that it is an "at least one" problem yet I am still struggling with it.
About 13% of the population is left handed. If 2 people are randomly selected, what is the probability that at least one is right handed?
Thanks!

Answer by stanbon(75887)

About 13% of the population is left handed. If 2 people are randomly selected, what is the probability that at least one is right handed?
----
Binomial Problem with n = 2 and p(left) = 0.13
----
P(1<= x <=2)
= 1 - P(x = 0)
= 1 - 0.87^3
= 1 - 0.6585
= 0.3415
===========================
Cheers,
Stan H.